A taxi driver travelling at constant speed of 14m/s runs past police check point. A police car parked on the road at this check point gives a chase at the instant the taxi driver passes the check point. Assuming that the police car can move at a constant acceleration of 3m/s2,

(a) how far from the check point does the police car catch up with the taxi driver?

(b) how long does it take?

(c) make a sketch of the velocity – time graph for both car on the same axes.


1
Expert's answer
2021-02-28T07:27:22-0500

a)


"s=vt=\frac{2v^2}{a}=\frac{2(14)^2}{3}=130.7\ m"

b)


"vt=0.5at^2\\t=\frac{2v}{a}=\frac{2(14)}{3}=9.33\ s"

c)